Grade 11Mechanics

A Plymouth with a mass of 2210 kg is moving along a straight stretch of road at 105 kmlh. It is followed by a Ford with mass 2080 kg moving at 43.5 kmlh. How fast is the center of mass of the two cars moving?

Simran Bhatia
11 Years agoGrade 11

1 Answer

Aditi Chauhan
11 Years ago

To find the velocity of the center of mass of the two cars, we use the formula for the velocity of the center of mass:

V_cm = (m₁v₁ + m₂v₂) / (m₁ + m₂)

where:
m₁ = 2210 kg (mass of Plymouth)
v₁ = 105 km/h (velocity of Plymouth)
m₂ = 2080 kg (mass of Ford)
v₂ = 43.5 km/h (velocity of Ford)

Step 1: Convert velocities to m/s
Since 1 km/h = 5/18 m/s, we convert the velocities:

v₁ = 105 × (5/18) = 29.167 m/s
v₂ = 43.5 × (5/18) = 12.083 m/s

Step 2: Compute the center of mass velocity
V_cm = (2210 × 29.167 + 2080 × 12.083) / (2210 + 2080)
= (644590.7 + 251326.4) / 4290
= 895917.1 / 4290
= 20.89 m/s

Step 3: Convert back to km/h
V_cm = 20.89 × (18/5) = 75.2 km/h

Final Answer:
The center of mass of the two cars is moving at 75.2 km/h.